# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ 3 Pseudodifferential operators acting on 1-forms
$\{\{Q_{\mathrm{prin}},R_{\mathrm{prin}}\}\}_{\alpha}{}^{\beta}:=\left(\frac{\partial[Q_{\mathrm{prin}}]_{\alpha}{}^{\kappa}}{\partial x^{\gamma}}-\Gamma^{\alpha^{\prime}}{}_{\gamma\alpha}[Q_{\mathrm{prin}}]_{\alpha^{\prime}}{}^{\kappa}+\Gamma^{\kappa}{}_{\gamma\kappa^{\prime}}[Q_{\mathrm{prin}}]_{\alpha}{}^{\kappa^{\prime}}\right)\frac{\partial[R_{\mathrm{prin}}]_{\kappa}{}^{\beta}}{\partial\xi_{\gamma}}\\ -\frac{\partial[Q_{\mathrm{prin}}]_{\alpha}{}^{\kappa}}{\partial\xi_{\gamma}}\left(\frac{\partial[R_{\mathrm{prin}}]_{\kappa}{}^{\beta}}{\partial x^{\gamma}}-\Gamma^{\kappa^{\prime}}{}_{\gamma\kappa}[R_{\mathrm{prin}}]_{\kappa^{\prime}}{}^{\beta}+\Gamma^{\beta}{}_{\gamma\beta^{\prime}}[R_{\mathrm{prin}}]_{\kappa}{}^{\beta^{\prime}}\right)$ (3.26)
@@
~$\widetilde{\nabla}_{\alpha}[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma}:=\frac{\partial[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma}}{\partial x^{\alpha}}-\Gamma^{\beta^{\prime}}{}_{\alpha\beta}[Q_{\mathrm{prin}}]_{\beta^{\prime}}{}^{\gamma}+\Gamma^{\gamma}{}_{\alpha\gamma^{\prime}}[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma^{\prime}},$
@@
It is possible to introduce genuine covariant derivatives on the cotangent bundle and reformulate formulae (3.6) and (3.26) in terms of the latter; see, for example, [16, Remark 4.5]. We refrain from doing this because it would take us away from the core subject of our paper.
@@
($\left.\left(\{\{Q_{\mathrm{prin}},R_{\mathrm{prin}}\}\}_{\alpha}{}^{\beta}\right)\right|_{x=z}=\left.\left(\widetilde{e}\,{}^{j}{}_{\alpha}\,\{(Q_{1/2})_{\mathrm{prin}},(R_{1/2})_{\mathrm{prin}}\}_{j}{}^{k}\,\widetilde{e}_{k}{}^{\beta}\right)\right|_{x=z}\,.$ (3.29)
@@
$[(Q_{1/2})_{\mathrm{prin}}]_{j}{}^{l}=\widetilde{e}_{j}{}^{\mu}\,[Q_{\mathrm{prin}}]_{\mu}{}^{\nu}\,\widetilde{e}\,{}^{l}{}_{\nu}\,,\qquad[(R_{1/2})_{\mathrm{prin}}]_{l}{}^{k}=\widetilde{e}_{l}{}^{\rho}\,[R_{\mathrm{prin}}]_{\rho}{}^{\sigma}\,\widetilde{e}\,{}^{k}{}_{\sigma}\,,$
@@ 4 Trace of pseudodifferential operators acting on 1-forms
$s<-d.$ (4.1)
Then [49, §12.1] $Q$ is an integral operator with continuous integral kernel,
@@
We introduced the factor $\rho(y)$ in the above integral in order to make the integral kernel $\mathfrak{q}_{j}{}^{k}(x,y)$ a (matrix-valued) scalar function on $M\times M$. Furthermore, if $Q$ is self-adjoint, then it is of trace class and we have
@@
where $(\operatorname{tr}\mathfrak{q})(x,y):=\mathfrak{q}_{j}{}^{j}(x,y)$ is the trace of the matrix-function $\mathfrak{q}_{j}{}^{k}(x,y)$. We define the scalar operator
@@
Of course, when condition (4.1) is not satisfied the integral kernel $\mathfrak{q}_{j}{}^{k}(x,y)$ appearing in formula (4.2) should be understood in the distributional sense (Schwartz kernel), but this does not prevent us from taking its matrix trace. The distribution $(\operatorname{tr}\mathfrak{q})(x,y)$ will be the Schwartz kernel of the scalar operator (4.4).
@@
acting on 1-forms. Here we encounter a problem: the quantity $\mathfrak{q}_{\alpha}{}^{\alpha}(x,y)$ is not a scalar on $M\times M$ because the tensor indices $\alpha$ and $\beta$ in $q_{\alpha}{}^{\beta}(x,y)$ ‘live’ at different points, $x$ and $y$ respectively. More precisely, the problem here is that the Schwartz kernel $\mathfrak{q}_{\alpha}{}^{\beta}(x,y)$ of the operator (4.5) is a two-point tensor: it is a covector at $x$ in the tensor index $\alpha$ and a vector at $y$ in the tensor index $\beta$. We overcome this impediment as follows.
@@
In what follows we will be raising and lowering indices in the 2-point tensor $Z_{\alpha}{}^{\beta}(x,y)$ using the Riemannian metric $g(x)$ in the first index and $g(y)$ in the second. Of course,
@@
$Z_{\beta}{}^{\alpha}(x,y)=Z^{\alpha}{}_{\beta}(y,x).$ (4.9)
Indeed, take the first identity (4.7), multiply it by $Z^{\alpha}{}_{\gamma}(x,y)$ and then use the second identity (4.8). This gives us $\,\delta^{\kappa}{}_{\gamma}\,Z_{\kappa}{}^{\beta}(y,x)\,=\,\delta_{\alpha}{}^{\beta}\,Z^{\alpha}{}_{\gamma}(x,y)\,$, which is equivalent to (4.9).
@@
$Z^{\alpha}{}_{\kappa}(x,\gamma(x,y;\tau))\,\mathfrak{q}_{\alpha}{}^{\beta}(x,y)\,Z_{\beta}{}^{\kappa}(y,\gamma(x,y;\tau))\,,$ (4.10)
where $\mathfrak{q}_{\alpha}{}^{\beta}(x,y)$ is the Schwartz kernel from (4.5) and $\tau\in[0,1]$. What happens here is that tensor indices are parallel transported to the common point $z=\gamma(x,y;\tau)$ on the geodesic $\gamma(x,y)$ connecting $x$ and $y$. Thus, the quantity (4.10) is a genuine scalar.
@@
$C^{\infty}(M\times M)\ni f(x,y)\mapsto Z^{\alpha}{}_{\kappa}(x,\gamma(x,y;\tau))\,Z_{\beta}{}^{\kappa}(y,\gamma(x,y;\tau))\,f(x,y)$
The quantity $\mathfrak{q}_{\alpha}{}^{\beta}(x,y)$ in (4.10) is a distribution (continuous complex-valued linear functional) acting on smooth $(1,1)$ two-point tensors. Formula (4.10) should be understood as a composition of the latter with the former.
@@
The proof of (4.17) is more subtle. Formula (3.1) tells us that, modulo an infinitely smooth contribution, the Schwartz kernel $\mathfrak{q}_{\alpha}{}^{\beta}(x,y)$ of the operator (4.5) reads
@@
where the symbol $q_{\alpha}{}^{\beta}(x,\xi)$ admits the asymptotic expansion (3.2). We also have
@@
Substituting (4.20) and (4.21) into (4.12) we arrive at a representation of the scalar pseudodifferential operator (4.13) with amplitude depending on $x$, $y$ and $\xi$. In order to write down the (left) symbol of the operator (4.13) we need to exclude the dependence on $y$, which is achieved in the standard manner, by means of integration by parts on the basis of the formula
$(x-y)^{\lambda}\,e^{i(x-y)^{\gamma}\xi_{\gamma}}=-i\,\frac{\partial e^{i(x-y)^{\gamma}\xi_{\gamma}}}{\partial\xi_{\lambda}}\,.$ (4.22)
This produces the following expressions for the leading and subleading terms of the symbol $(\operatorname{\mathfrak{tr}}q)(x,\xi)$ of the scalar pseudodifferential operator (4.13):
@@ 5 The projection operators $P_{0}$ and $P_{\pm}$
and we write $Q=R\mod\Psi^{-\infty}$ if $Q-R$ is an integral operator with (infinitely) smooth integral kernel.
@@
Let $f_{j}$ be the orthonormalised eigenfunctions corresponding to the eigenvalues (2.17) of $\,-\Delta\,$. Put
@@
$P_{0}:=\sum_{j=1}^{+\infty}v_{j}\langle v_{j}\,,\,\cdot\,\rangle.$ (5.3)
@@
$\Delta^{-1}:=-\sum_{j=1}^{+\infty}\mu_{j}^{-1}f_{j}\langle f_{j}\,,\,\cdot\,\rangle$ (5.4)
@@
$P_{-}=\sum_{j=-1}^{-\infty}u_{j}\langle u_{j}\,,\,\cdot\,\rangle.$
@@
$\bm{\Delta}:=-(\mathrm{d}\delta+\delta\mathrm{d})=-\sum_{j=1}^{+\infty}\mu_{j}\,v_{j}\langle v_{j}\,,\,\cdot\,\rangle-\sum_{j\in\mathbb{Z}\setminus\{0\}}\lambda_{j}^{2}\,u_{j}\langle u_{j}\,,\,\cdot\,\rangle$ (5.8)
@@
$(-\bm{\Delta})^{s}:=\sum_{j=1}^{+\infty}\mu_{j}^{s}\,v_{j}\langle v_{j}\,,\,\cdot\,\rangle+\sum_{j\in\mathbb{Z}\setminus\{0\}}|\lambda_{j}|^{2s}\,u_{j}\langle u_{j}\,,\,\cdot\,\rangle$ (5.9)
@@
$[\operatorname{Id}_{\mathrm{prin}}]_{\alpha}{}^{\beta}=\delta_{\alpha}{}^{\beta}\,,$ (5.12)
$[-\bm{\Delta}_{\mathrm{prin}}]_{\alpha}{}^{\beta}=\|\xi\|^{2}\,\delta_{\alpha}{}^{\beta}\,.$ (5.13)
@@
$P_{\aleph}^{*}=P_{\aleph},\qquad P_{\aleph}P_{\beth}=\delta_{\aleph\beth}P_{\aleph}\,,\qquad\sum_{\aleph}\ P_{\aleph}=\mathrm{Id}\,,\qquad[\operatorname{curl},P_{\aleph}]=0\,,$ (5.14)
@@
$[\operatorname{curl},\widetilde{P}_{\aleph}]=0\mod\Psi^{-\infty}\,.$ (5.15d)
@@
The claim follows from a straightforward adaptation of [17, Theorem 2.2] to the setting of the current paper. The original result from [17] was formulated under the assumption of ellipticity but ellipticity was never used in establishing the existence of projections. What was used is simplicity of eigenvalues of the principal symbol, which holds for $\operatorname{curl}$. Furthermore, the original result from [17] was formulated under the assumption that the operator acts on columns of half-densities, but this assumption was never used in an essential way; what was used is the presence of an inner product. ∎
@@
The proof is analogous to that of Theorem 5.1, only additionally taking into account [17, Theorem 4.1]. ∎
@@
Let introduce refined notation for the principal symbol. Namely, we denote by $(\,\cdot\,)_{\mathrm{prin},s}$ the principal symbol of the expression within brackets, regarded as an operator in $\Psi^{-s}$. To appreciate the need for such notation, consider the following example. Let $Q$ and $R$ be pseudodifferential operators in $\Psi^{-s}$ with the same principal symbol. Then, as an operator in $\Psi^{-s}$, $Q-R$ has vanishing principal symbol: $(Q-R)_{\mathrm{prin},s}=0$. But this tells us that $Q-R$ is, effectively, an operator in $\Psi^{-s-1}$ and, as such, it may have nonvanishing principal symbol $(Q-R)_{\mathrm{prin},s+1}$. This refined notation will be used whenever there is risk of confusion.
@@
The operators $P_{0}$, $P_{\pm}$ can be explicitly constructed, modulo $\Psi^{-\infty}$, as follows.
@@
The algorithm presented here is a straightforward adaptation of that from [17, subsection 4.3], which can be applied to the operator $\operatorname{curl}$ in view of Theorem 5.1. ∎
@@
The paper [17] offers two different algorithms for the construction of an orthonormal basis of pseudodifferential projections. The first algorithm, using the full set of conditions from Theorem 5.1, is given in [17, subsection 3.4]. The second algorithm, using the minimal set of conditions from Theorem 5.2, is given in [17, subsection 4.3]. Of course, the end result is the same, modulo $\Psi^{-\infty}$, but the latter is much simpler.
@@
$[(X_{\aleph,1})_{\mathrm{prin}}(z,\zeta)]_{\alpha}{}^{\beta}=0$ (5.22)
@@
$[R_{\aleph,1}]_{\alpha}{}^{\beta}(z,\zeta)=-[(P_{\aleph}^{2})_{\mathrm{sub}}]_{\alpha}{}^{\beta}(z,\zeta)=-\frac{i}{2}\{\{(P_{\aleph})_{\mathrm{prin}}\,,(P_{\aleph})_{\mathrm{prin}}\}\}_{\alpha}{}^{\beta}(z,\zeta)=0\,.$ (5.23)
@@
Formula (5.23) immediately implies $[S_{\aleph,k}]_{\alpha}{}^{\beta}(z,\zeta)=0$. Hence, on account of Lemma 3.6, we have
$[T_{\aleph,1}]_{\alpha}{}^{\beta}(z,\zeta)=\frac{i}{2}\{\{(P_{\aleph})_{\mathrm{prin}}\,,\operatorname{curl}_{\mathrm{prin}}\}\}(z,\zeta)-\frac{i}{2}\{\{\operatorname{curl}_{\mathrm{prin}}\,,(P_{\aleph})_{\mathrm{prin}}\}\}(z,\zeta)\,.$
@@
Observe that (1.8)–(1.11) and (3.2) give us explicit formulae for the homogeneous components of degree $-1$ of the symbols of the operators $P_{0}$ and $P_{\pm}$ in arbitrary local coordinates.

### v4

@@ 3 Pseudodifferential operators acting on 1-forms
[…]
@@
$\widetilde{\nabla}_{\alpha}[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma}:=\frac{\partial[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma}}{\partial x^{\alpha}}-\Gamma^{\beta^{\prime}}{}_{\alpha\beta}[Q_{\mathrm{prin}}]_{\beta^{\prime}}{}^{\gamma}+\Gamma^{\gamma}{}_{\alpha\gamma^{\prime}}[Q_{\mathrm{prin}}]_{\beta}{}^{\gamma^{\prime}}$,
@@
It is possible to introduce genuine covariant derivatives on the cotangent bundle and reformulate formulae (3.6) and (3.26) in terms of the latter; see, for example, [17, Remark 4.5]. We refrain from doing this because it would take us away from the core subject of our paper.
@@
$\left.\left(\{\{Q_{\mathrm{prin}},R_{\mathrm{prin}}\}\}_{\alpha}{}^{\beta}\right)\right|_{x=z}=\left.\left(\widetilde{e}\,{}^{j}{}_{\alpha}\,\{(Q_{1/2})_{\mathrm{prin}},(R_{1/2})_{\mathrm{prin}}\}_{j}{}^{k}\,\widetilde{e}_{k}{}^{\beta}\right)\right|_{x=z}\,.$ (3.29)
@@
$(Q_{1/2})_{\mathrm{prin}}]_{j}{}^{l}=\widetilde{e}_{j}{}^{\mu}\,[Q_{\mathrm{prin}}]_{\mu}{}^{\nu}\,\widetilde{e}\,{}^{l}{}_{\nu}\,,\qquad[(R_{1/2})_{\mathrm{prin}}]_{l}{}^{k}=\widetilde{e}_{l}{}^{\rho}\,[R_{\mathrm{prin}}]_{\rho}{}^{\sigma}\,\widetilde{e}\,{}^{k}{}_{\sigma}\,$,
@@ 4 Trace of pseudodifferential operators acting on 1-forms
s<d.s<-d. (4.1)
Then [59, §12.1] $Q$ is an integral operator with continuous integral kernel,
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$C^{\infty}(M\times M)\ni f(x,y)\mapsto Z^{\alpha}{}_{\kappa}(x,\gamma(x,y;\tau))\,Z_{\beta}{}^{\kappa}(y,\gamma(x,y;\tau))\,f(x,y)\,.$ (4.11)
[…]
@@
[…]
@@
[…]
@@
Substituting (4.21) and (4.22) into (4.13) we arrive at a representation of the scalar pseudodifferential operator (4.14) with amplitude depending on $x$, $y$ and $\xi$. In order to write down the (left) symbol of the operator (4.14) we need to exclude the dependence on $y$, which is achieved in the standard manner — recall (3.13). This produces the following expressions for the leading and subleading terms of the symbol $(\operatorname{\mathfrak{tr}}q)(x,\xi)$ of the scalar pseudodifferential operator (4.14):
@@ 5 The projection operators $P_{0}$ and $P_{\pm}$
and we write $Q=R\mod\Psi^{-\infty}$ if QRQ-R is an integral operator with (infinitely) smooth integral kernel.
@@
Let $f_{j}$ be the orthonormalised eigenfunctions corresponding to the eigenvalues (B.12) of $\Delta\,$. Put
@@
$P_{0}:=\sum_{j=1}^{+\infty}v_{j}\langle v_{j}\,$$\cdot\,\rangle$. (5.3)
@@
$\Delta^{-1}:=-\sum_{j=1}^{+\infty}\mu_{j}^{-1}f_{j}\langle f_{j}\,$$\cdot\,\rangle$ (5.4)
@@
$P_{-}=\sum_{j=-1}^{-\infty}u_{j}\langle u_{j}\,$$\cdot\,\rangle$.
@@
$\bm{\Delta}:=-(\mathrm{d}\delta+\delta\mathrm{d})=-\sum_{j=1}^{+\infty}\mu_{j}\,v_{j}\langle v_{j}\,,\,\cdot\,\rangle-\sum_{j\in\mathbb{Z}\setminus\{0\}}\lambda_{j}^{2}\,u_{j}\langle u_{j}\,$$\cdot\,\rangle$ (5.8)
@@
$(-\bm{\Delta})^{s}:=\sum_{j=1}^{+\infty}\mu_{j}^{s}\,v_{j}\langle v_{j}\,,\,\cdot\,\rangle+\sum_{j\in\mathbb{Z}\setminus\{0\}}|\lambda_{j}|^{2s}\,u_{j}\langle u_{j}\,$$\cdot\,\rangle$ (5.9)
@@
$\operatorname{Id}_{\mathrm{prin}}]_{\alpha}{}^{\beta}=\delta_{\alpha}{}^{\beta}\,$, (5.12)
$\bm{\Delta}_{\mathrm{prin}}]_{\alpha}{}^{\beta}=\|\xi\|^{2}\,\delta_{\alpha}{}^{\beta}\,$. (5.13)
@@
$P_{\aleph}^{*}=P_{\aleph},\qquad P_{\aleph}P_{\beth}=\delta_{\aleph\beth}P_{\aleph}\,,\qquad\sum_{\aleph}\ P_{\aleph}=\mathrm{Id}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mod\Psi^{-\infty}}\,,\qquad[\operatorname{curl},P_{\aleph}]=0\,$, (5.14)
@@
$\operatorname{curl},\widetilde{P}_{\aleph}]=0\mod\Psi^{-\infty}\,.$ (5.15d)
@@
The claim follows from a straightforward adaptation of [18, Theorem 2.2] to the setting of the current paper. The original result from [18] was formulated under the assumption of ellipticity but ellipticity was never used in establishing the existence of projections. What was used is simplicity of eigenvalues of the principal symbol, which holds for $\operatorname{curl}$. Furthermore, the original result from [18] was formulated under the assumption that the operator acts on columns of half-densities, but this assumption was never used in an essential way; what was used is the presence of an inner product. ∎
@@
The proof is analogous to that of Theorem 5.1, only additionally taking into account [18, Theorem 4.1]. ∎
@@
Let introduce refined notation for the principal symbol. Namely, we denote by $(\,\cdot\,)_{\mathrm{prin},s}$ the principal symbol of the expression within brackets, regarded as an operator in $\Psi^{-s}$. To appreciate the need for such notation, consider the following example. Let $Q$ and $R$ be pseudodifferential operators in $\Psi^{-s}$ with the same principal symbol. Then, as an operator in $\Psi^{-s}$, QRQ-R has vanishing principal symbol: $(Q-R)_{\mathrm{prin},s}=0$. But this tells us that QRQ-R is, effectively, an operator in $\Psi^{-s-1}$ and, as such, it may have nonvanishing principal symbol $(Q-R)_{\mathrm{prin},s+1}$. This refined notation will be used whenever there is risk of confusion.
@@
The operators $_{0}, P±P_{\pm}$ can be explicitly constructed, modulo $\Psi^{-\infty}$, as follows.
@@
    Recall that $h^{(\,\cdot\,)}$ is defined in accordance with (2.3) and (1.7).
@@
The algorithm presented here is a straightforward adaptation of that from [18, subsection 4.3], which can be applied to the operator $\operatorname{curl}$ in view of Theorem 5.1. ∎
@@
The paper [18] offers two different algorithms for the construction of an orthonormal basis of pseudodifferential projections. The first algorithm, using the full set of conditions from Theorem 5.1, is given in [18, subsection 3.4]. The second algorithm, using the minimal set of conditions from Theorem 5.2, is given in [18, subsection 4.3]. Of course, the end result is the same, modulo $\Psi^{-\infty}$, but the latter is much simpler.
@@
$(X_{\aleph,1})_{\mathrm{prin}}(z,\zeta)]_{\alpha}{}^{\beta}=0$ (5.22)
@@
$R_{\aleph,1}]_{\alpha}{}^{\beta}(z,\zeta)=-[(P_{\aleph}^{2})_{\mathrm{sub}}]_{\alpha}{}^{\beta}(z,\zeta)=-\frac{i}{2}\{\{(P_{\aleph})_{\mathrm{prin}}\,,(P_{\aleph})_{\mathrm{prin}}\}\}_{\alpha}{}^{\beta}(z,\zeta)=0\,.$ (5.23)
@@
Formula (5.23) immediately implies $S_{\aleph,k}]_{\alpha}{}^{\beta}(z,\zeta)=0$. Hence, on account of Lemma 3.6, we have
$T_{\aleph,1}]_{\alpha}{}^{\beta}(z,\zeta)=\frac{i}{2}\{\{(P_{\aleph})_{\mathrm{prin}}\,,\operatorname{curl}_{\mathrm{prin}}\}\}(z,\zeta)-\frac{i}{2}\{\{\operatorname{curl}_{\mathrm{prin}}\,,(P_{\aleph})_{\mathrm{prin}}\}\}(z,\zeta)\,.$
@@
Observe that (1.8)–(1.11) and (3.2) give us explicit formulae for the homogeneous components of degree 1-1 of the symbols of the operators $P_{0}$ and $P_{\pm}$ in arbitrary local coordinates.